The width of a rectangle is three units less than the length. If the area is 28 square units, then find the dimensions of the rectangle.
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

width=x
length=x+3
x(x+3)=28
x^2+3x-28=0
(x+7)(x-4)=0
x=4, only positive root
x+3=7
4 x 7 units

Answer by ikleyn(52844)

.
7 by 4 units.

7 and 4 are the only positive (integer) factors of 28 that differ by 3.

So, this problem is for mental solution.

But if you want to see how to solve it using equation, look into the lesson 

    - Problems on the area and the dimensions of a rectangle

in this site.


With equations, the solution is THIS:

Area A = LW (the product of the length L and the width W).

From the condition, W = L-3. Substitute it into the equation for area.
You will get the quadratic equation

L*(L-3) = 28,

L^2 - 3L - 28 = 0.

Factor left side:

(L-7)*(L+4) = 0.

The only positive root is L=3, which gives you the answer: L = 7 units, W = 7-3 = 4 units.

Solved.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic
"Dimensions and the area of rectangles and circles and their elements".